Massachusetts Institute of Technology Handout 20 6.042J/18.062J: Mathematics for Computer Science October 20, 2000 Professors Nancy Lynch and Srinivas Devadas

Mini-Quiz 6

1. Write your name:

2. (modified from Rosen, Section 1.8, Problem 13) Show that 2^{n} is O(4^{n}) but that 4^{n} is not O(2^{n}).